In a jar, there are 45 nickels. The number of pennies and nickels together equal the number of dimes. There are 2 quarters for every 3 nickels and 2 dimes for every quarter.
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

pennies=x
dimes=x+45
2 quarters for 3 nickels means 30 quarters
2 dimes for every quarter means 60 dimes
pennies has to be 15

Answer by ikleyn(52925)

.
1.  Since "there are 2 quarters for every 3 nickels", you can easily determine the number of quarters.

    For it, divide 45 by 3 to get 15 and then multiply by 2 to get the number of quarters: 30.



2.  Since "There are . . . 2 dimes for every quarter", the number of dimes is 2*30 = 60.



3.  #Pennies + #Nickels = #Dimes = 60  ====>

    #Pennies + 45                = 60  ====> #Pennies = 60 - 45 = 15.



Answer.  15 pennies,  45 nickels (given),  60  dimes  and  30 quarters.